# Optimizing Multi-level Magic State Factories for Fault-Tolerant Quantum Architectures

### v1

@@ Abstract
We propose a novel technique for optimizing a modular fault-tolerant quantum computing architecture, taking into account any desired space–time trade-offs between the number of physical qubits and the fault-tolerant execution time of a quantum algorithm. We consider a concept architecture comprising a dedicated zone as a multi-level magic state factory and a core processor for efficient logical operations, forming a supply chain network for production and consumption of magic states. Using a heuristic algorithm, we solve the multi-objective optimization problem of minimizing space and time subject to a user-defined error budget for the success of the computation, taking the performance of various fault-tolerant protocols such as quantum memory, state preparation, magic state distillation, code growth, and logical operations into account. As an application, we show that physical quantum resource estimation reduces to a simple model involving a small number of key parameters, namely, the circuit volume, the error prefactors ($\mu$) and error suppression rates ($\Lambda$) of the fault-tolerant protocols, and an allowed slowdown factor ($\beta$). We show that, in the proposed architecture, $10^{5}$$10^{8}$ physical qubits are required for quantum algorithms with $T$-counts in the range $10^{6}$$10^{15}$ and logical qubit counts in the range $10^{2}$$10^{4}$, when run on quantum computers with quantum memory $\Lambda$ in the range 3–10, for all slowdown factors $\beta\geq 0.2$.
@@ I Introduction
Fault-tolerant compilation of quantum algorithms is especially more complicated than that of classical computer programs because the final physical circuit depends on the specific noise characteristics of the quantum processing unit (QPU). It is commonly understood that the number of non-Clifford logical operations (i.e., the $T$-count of the algorithm) is a good indicator of the approximate cost of running the quantum algorithm. However, assembling the quantum program to run in hours, days, or even months of exact physical circuits on a quantum computer with millions of qubits is much more complicated. Recently, several physical resource estimation platforms have been developed [1, 2] to provide a finer analysis of fault-tolerant quantum computation (FTQC) at utility scale. However, these resource estimators rely on involved bookkeeping of physical qubit counts for various procedures, and therefore are not easily adaptable to alternative circuit decompositions (i.e., beyond Clifford+$T$, e.g., by including Toffoli gates), physical architectures (e.g., for using different state distillation and injection protocols), or the various and possibly very complicated descriptions of the noise of the hardware. This paper is intended to provide a systematic and simple approach to FTQC architecture design, analysis, and optimization.
@@
We present a framework for constructing flexible varieties of FTQC architectures in a modular fashion by allocating dedicated zones across a 2D physical qubit layout for the execution of specific fault-tolerant protocols. To this end, we view FTQC as the continual production and consumption of various types of expensive quantum resources, such as the production of lower-fidelity logical |$\ket{T}$ or |$\ket{CCZ}$ states in one zone and their consumption in another zone for obtaining higher-fidelity magic states via a distillation protocol [3]. We consider rotated surface codes as the typical choice of quantum error correction (QEC) codes, and rely on lattice surgery for the execution of entangling gates and seamless routing and teleportation of quantum information across a quantum bus [4, 5].
At utility scale, it is economical to dedicate small numbers of physical code patches as buffer registers connecting the FTQC architecture zones to each other. Newly produced resources in one zone are placed in a buffer, where they are fault-tolerantly maintained via continued stabilizer measurements and decoding (quantum memory) for consumption in a subsequent zone. One or many terminal core processors consume the final highest-fidelity resources to perform the logical gates of the quantum program. We refer the reader to Section II and Figs. 2 and 3 for further details.
(a) Space-time costs for varying quantum circuits
(b) Space-time costs for varying quantum hardware
Figure 1: Space and time cost estimates of an efficient fault-tolerant quantum architecture. (a) Estimates for combinations of $T$-count $T=10^{6}$ to $10^{15}$ and logical data qubits $Q=10^{2}$ to $10^{4}$ in a quantum circuit, based on an FTQC model with a quantum memory error prefactor of $\mu_{\text{mem}}=3.8\times 10^{-3}$ and an error suppression rate of $\Lambda_{\text{mem}}=10$, and a magic state preparation protocol with a linear form with $\mu_{\text{prep}}=1.44\times 10^{-4}$ and $\Lambda_{\text{prep}}=2.5\times 10^{-5}$. (b) Estimates for combinations of $\Lambda_{\text{mem}}=3$ to $10$ and a magic state preparation protocol with $\Lambda_{\text{prep}}=0$ and $\mu_{\text{prep}}=e_{\text{prep}}=10^{-2}$ to $10^{-8}$ for a circuit with $T=10^{12}$ and $Q=10^{2}$. Each data point corresponds to a slowdown factor $\beta$, relative to a serial execution ($\beta=1$) of the quantum circuit. A value of $\beta<1$ reflects a scenario where the magic state factory (MSF) is expanded to meet the magic state consumption rate when parallelizing non-Clifford gates, resulting in faster execution. Conversely, $\beta>1$ indicates a reduction in MSF size, leading to a longer runtime. While runtime is heavily influenced by the number of non-Clifford gates in the circuit and the MSF size, our estimates demonstrate that quantum computers require between $10^{5}$ and $10^{8}$ physical qubits across all scenarios evaluated.
The buffer registers provide several advantages. Practically speaking, they allow for scheduling operations within each zone independently and asynchronously from the other zones. Moreover, since fault-tolerant protocols typically fail with certain non-zero probability, keeping the buffers fully stacked with their associated resource states guarantees that the operations of any dependent zones do not halt since the probability of encountering a completely empty buffer is extremely small. More generally, the oversupply and undersupply of resource states in a buffer can be intentionally used to adjust the space–time trade-offs for FTQC. For example, a magic state factory (MSF) may require many magic state distillation units to keep its buffer fully stacked, while it is possible for a smaller number of distillation units to still be able to fault-tolerantly execute the quantum algorithm if some slowdown caused by occasionally empty buffer registers can be tolerated.
The key idea in this work is to determine the size of each FTQC architecture zone by choosing the number of FTQC protocol units (e.g., 15:1 magic state distillation units, or the like) in that zone in such a way that the supply and demand of the interconnected zones are balanced. This balance, however, can be broken, for the assembly of the program using fewer physical qubits by incurring a longer execution time. However, reducing the (logical) size of the MSF increases the idle time in the core processor, which increases the overall accumulated error rates, potentially requiring larger code distances. Our study provides a detailed analysis of such trade-offs. We model the optimization of space and time as a bi-objective optimization problem in Section III. By solving this optimization problem, we observe that the space and time costs of any FTQC can be efficiently predicted from a small set of key attributes of the quantum program and the quantum computer executing it, and the potential space–time trade-offs for the architecture can be identified.
For clarity we distinguish FTQC compilation and assembly as follows. We assume that the quantum program is given in an intermediate representation (IR) language, also called assembly, after compilation from a higher-level programming environment. Examples of common quantum IR languages are QASM [6] and QIR [7]. However, in this paper, we focus on translation at a lower level from assembly to machine instructions (i.e., QEC rounds and decoder instructions) performed by an assembler. Different IR programs may be transpiled into each other efficiently [5, 4]. Therefore, we assume Pauli-based lattice surgery as our IR of choice. For the assembler, the two relevant attributes of the FTQC program are denoted by the Greek small letters $\alpha$ and $\beta$, representing the average size of lattice surgeries in the core processor relative to the number of logical qubits and a user-defined slowdown factor within the buffers, respectively. Our empirical studies showed that $\alpha$ is not a significant factor for the assembly process, so we may omit it to obtain an even simpler model depending on only the slowdown factor $\beta$. The noise profile of the QPU is also important for the assembler. We observe that this information can be reduced to a small number of parameters for a predictive model of the logical error rates of each FTQC protocol, namely $\mu$ and $\Lambda$, respectively representing the logical error prefactor and suppression rate of the protocol. Prior literature has used $\Lambda=\Lambda_{\text{mem}}$ to quantify the performance of the quantum hardware in a fault-tolerant quantum memory experiment on a single surface code patch [8]. $\Lambda_{\text{mem}}$ describes the asymptotic multiple of improvement in the logical fidelity of the QEC code when increasing its distance by 2. More comprehensively, a predictive model for the logical error rate of every fault-tolerant protocol can be devised using additional model parameters. Figure 1 provides a summary of the resource estimates for quantum circuits with varying $T$-counts, logical qubit counts $Q$, and quantum computers with varying $\Lambda_{\text{mem}}$ and first-stage preparation error rate $e_{\text{prep}}$.
Overall, the main contributions of our paper are the following:
  • A new modular architecture design procedure for FTQC with at least a multi-level MSF and a core processor;
  • An optimization framework that decides the size required for the MSF and the code distances of all logical qubits while minimizing space and time costs; and
  • A sensitivity analysis of space and time trade-offs for different quantum circuits and quantum computers.
The paper is organized as follows. Section II introduces our proposed architecture. Section III introduces the error models driven from numerical and experimental hardware noise data. Section IV formalizes the bi-objective optimization problem to be solved and the optimization framework we used to solve the problem. Section V presents our numerical results for trading resource estimations on varying quantum circuits and quantum hardware noise profiles. Finally, Section VI offers concluding remarks.
@@ II The Fault-Tolerant Architecture
We assume rotated surface codes as our QEC scheme of choice and assume that the program to be run is transpiled to multi-qubit Pauli rotations of the angles $\pi$, $\pi/2$, and $\pi/4$ for Clifford operations, and $\pi/8$ for non-Clifford operations [9, 10]. These Pauli rotations are in turn implemented using fault-tolerant lattice surgery [5, 11, 4]. A core processor is designed to sequentially perform multi-qubit, long-range entanglements for sets of $\pi/8$ Pauli rotation measurements using auxiliary qubits to connect all the logical qubits required by a given logical operation. The core (see Section II.1) is connected to an external multi-level MSF as detailed in Section II.2 that supplies high-fidelity magic states required for the implementation of these logical operations.
@@ II.1 The Core Processor
The core processor is the central component of the architecture, where logical operations are performed on the logical qubits. It comprises two key elements: the memory fabric and a buffer. The memory fabric is responsible for storing the quantum information required by the computation and performing its logical operations. We follow the fast block layout of Ref. [5], where logical qubits are arranged in two-tile, two-qubit data patches connected to a quantum bus using a square arrangement for performing lattice surgery. This is a time-efficient layout as all Pauli operators—$X$, $Z$, and $Y$ being the combination of both—are directly accessible by the quantum bus. In an exact square arrangement, the $Q$ logical qubits requested in the algorithm would require a space of $12Q+\sqrt{8Q}+$ 1 tiles plus the tiles required for the buffer space. To save on space, we shorten the last column when $\sqrt{Q/2}$ is not an integer.
The buffer stores the resources prepared in the external MSF and holds them until they are consumed by the $\pi/8$ rotations. The buffer is used to simplify access to magic states so that the memory fabric does not have to directly interface with the MSF. It allows the execution of auto-corrected $\pi/8$ rotations in a single time step, while receiving and preparing new magic states in parallel following the protocol suggested in Ref. [5]. It includes spaces dedicated for magic state growth and storage, which are used for expanding magic states to larger code distances and preparing them for consumption, respectively.
An example layout of the core processor is shown in Fig. 2  for a circuit requiring 18 logical qubits to store its data. The example shows multiple storage and growth units to provide robustness against fluctuations in demand, resulting from parallelization of $\pi/8$ rotations, and production, resulting from distillation failures. More connection points between the memory fabric and the buffer can be added to allow greater parallelization of operations. We note that if the production rate of magic states in the MSF is slower than the consumption rate of magic states in the memory fabric, extra storage capacity is not required; therefore, the buffer space can be reduced in size.
@@
Figure 2: Example layout of the core processor. The memory fabric is composed of $Q=18$ logical qubits distributed in nine two-tile, two-qubit data patches (shown in purple) surrounded by bus qubits (green), while the buffer is composed of magic state storage patches (yellow) dedicated to storing magic states produced in the MSF. This storage space is designed to allow auto-corrected $\pi/8$ rotations by connecting the storage qubits to an ancillary qubit (pink) using lattice surgery in parallel to the lattice surgery used to connect the data patches to the magic state consumed. Magic state growth units are also presented to allow magic state expansion if magic states produced are of a different size than is required.
@@ II.2 Magic State Factory
The MSF is a module of the fault-tolerant architecture dedicated to the production of high-fidelity magic states required to implement the $\pi/8$ rotation gates. Magic states are initially prepared using physical operations in a magic state preparation unit. As these are highly error-prone states that most likely have fidelities below those required for fault tolerance, rounds of magic state distillation are performed using magic state distillation units, where the magic states at one particular level are consumed to create higher-fidelity magic states. Once ready, the magic states distilled are expanded to a larger code distance, if required, in magic state growth units, and transferred to a distillation unit in the next level. When magic states reach the required fidelity, the distillation process has been completed and the magic states created at the highest distillation level are sent to the core processor to be consumed when performing the $\pi/8$ rotation gates. Figure 3 shows an example of the architecture described.
Figure 3: Architecture of the MSF and its interface with the core processor buffer. Magic states are first prepared using dedicated preparation units following a preparation protocol. They are then dispatched to a buffer where they remain until being requested by its corresponding distillation unit. The higher-fidelity magic states are prepared in the distilling ports of these units. Once ready, they proceed to growth units dedicated to a next-level distillation unit. The process repeats until the highest level sends a magic state to the core processor. In this example, two distillation levels are used, composed of four and two distillation units each from lowest to highest.
@@
As a good example that is used as a testbed throughout this paper, the 15:1 distillation protocol [13] uses 15 lower-fidelity magic states and produces a single higher-fidelity magic state. An analysis of this distillation protocol shows that if the input error rates of the magic states are $e_{\text{in}}$ and the logical error rates of the individual logical gates are $e_{\text{mem}}$, then the output error is bounded by $O(e_{\text{in}}^{3}+e_{\text{mem}})$ from the distillation protocol. From this, one can determine the output fidelity of a single magic state distillation unit given the input fidelities and code distances. Additionally, it is possible to approximate the probability of a successful distillation for this protocol using numerical simulations [11]. Figure 3 shows an architecture for the 15:1 protocol customized to run in 13 logical operations, accounting for 11 gates, one cycle for correction, and one cycle for emptying the distillation port, at a given code distance. This architecture requires at least 28 surface code tiles, much more than the five logical qubits required for the protocol itself.
@@
The very first level of the multi-level MSF consumes logical magic states prepared from physical non-Clifford states that are fault-tolerantly grown to the logical state of a full surface code patch. These protocols generally involve the creation of a physically relevant state (such as a physical magic state), and then the injection of this magic state into a larger quantum error-correcting code [16, 17]. The result is a magic state of fidelity $e_{\text{prep}}$ at a specified code distance. Different protocols require different physical operations, and have different time scales, but most require $\sim$1–2 logical cycles, where a logical cycle requires $d$ rounds of QEC at the final code distance $d$. The physical execution of the protocols on a 2D layout causes additional error to be accumulated which can also be numerically estimated. At the end of the process, we obtain the error rate $e_{\text{in}}$ for the initial level of magic state distillation.
@@
Although distillation levels can be encoded using the same code distances, it is more space efficient to use different encodings for the logical qubits at different levels, usually in increasing order. Therefore, a growth protocol must be used between the MSF levels to increase the code distance of the magic states to match the distance of the next distillation level or that of the core processor. These growth protocols operate similarly to the preparation protocols, in which a state encoded in a smaller state is surrounded by additional qubits in specific states, followed by QEC rounds in the grown patch. These growth protocols generally have a small probability of failure, and these probabilities can nevertheless be numerically estimated.
@@
As magic state distillation is a probabilistic procedure, and thus each distillation unit has some chance of not outputting a state, our proposed architecture also includes a buffer register where magic states can remain idle while waiting to be consumed. In this manner, we can reduce the probability of no magic state existing when one is needed. These zones essentially need only to be a single logical tile where the magic state can sit, but having access to the tiles to both load and unload the magic states means they should generally be lined up. Additionally, when magic states interact with the memory fabric, we typically also need access to specific ancillary tiles to facilitate the logical operations that utilize magic states, so these additional ancillary regions are located near the buffer registers to decrease the required logical runtimes of operations involving magic states (see Fig. 3).
III Hardware Error Modelling

### v2

@@ Abstract
We propose a novel technique for optimizing a modular fault-tolerant quantum computing architecture, taking into account any desired space–time trade-offs between the number of physical qubits and the fault-tolerant execution time of a quantum algorithm. We consider a concept architecture comprising a dedicated zone as a multi-level magic state factory and a core processor for efficient logical operations, forming a supply chain network for production and consumption of magic states. Using a heuristic algorithm, we solve the multi-objective optimization problem of minimizing space and time subject to a user-defined error budget for the success of the computation, taking the performance of various fault-tolerant protocols into account. As an application, we show that physical quantum resource estimation reduces to a simple model involving a small number of key parameters, namely, the circuit volume, the error prefactors ($\mu$) and error suppression rates ($\Lambda$) of the fault-tolerant protocols, the reaction time ($\gamma$), and an allowed slowdown factor ($\beta$).
@@ I Introduction
Fault-tolerant compilation of quantum algorithms is especially more complicated than that of classical computer programs because the final physical circuit depends on the specific noise characteristics of a quantum processing unit (QPU). This information is required to decide the types and distances of error correcting codes used, and to thereby estimate the fidelities of various fault-tolerant protocols performed during the computation. Since utility-scale computations may require $10^{6}$$10^{12}$ non-Clifford operations performed during several days or even months on complex and expensive fault-tolerant quantum computing (FTQC) stacks, it is crucial to minimize the overhead of fault tolerance while bounding the accumulated errors of logical operations with high confidence.
In this paper, we provide a systematic approach to optimizing the design and sizes of various zones of a modular FTQC architecture. We call this procedure the assembly of an FTQC program, which has been previously compiled from a higher-level programming environment to a quantum assembly or intermediate-representation (IR) language (such as QASM [1] and QIR [2]). The key idea is to determine the size of each FTQC architecture zone by choosing the number of FTQC protocol units (e.g., 15:1 magic state distillation units, or the like) in that zone in such a way that the supply and demand of quantum states between the interconnected zones are balanced. This balance can be broken if a smaller total number of physical qubits or a faster execution time is desired.
@@
We first present a framework for constructing flexible varieties of FTQC architectures in a modular fashion by allocating dedicated zones across a 2D physical qubit layout for the execution of specific fault-tolerant protocols. To this end, we view FTQC as the continual production and consumption of various types of expensive quantum resources, such as the production of lower-fidelity logical |$\ket{T}$ or |$\ket{CCZ}$ states in one zone and their consumption in another zone for obtaining higher-fidelity magic states via a distillation protocol [3]. We use rotated surface codes as the typical choice of quantum error correction (QEC) codes, and rely on lattice surgery for the execution of entangling gates and routing of quantum information across a quantum bus [4, 5]. We also introduce buffer registers between the zones to faciliate asynchronous operation of interconnected zones. Such asynchronous operations are caused by differences in logical cycle times across the architecture and the stochastic nature of magic state distillation. The buffer registers ensure that the quantum computer operates in a long-term steady state, and reduce the chance of execution-time halting. Newly produced resources in one zone are stored and fault-tolerantly maintained via continued stabilizer measurements and decoding (quantum memory) in the buffers until their consumption in a subsequent zone. One or many terminal core processors consume the final highest-fidelity resources to perform the logical gates of the quantum program. We refer the reader to Section II and Figs. 1 and 2 for further details on the proposed architecture.
Our process shines particularly well when an MSF consists of heterogeous zones with various code distances or distillation protocols. In contrast to previous designs such as that of Ref. [6], which reuse the same area in the MSF for sequential distillation, using dedicated zones for distillation reduces the number of distillation units required in the MSF since less time is required for distillation. Dedicated zones also allow failed protocols to immediately and asynchronously restart, increasing the utilization of resources (see Appendix C for further comparative discussions). Our assembly method can optimize the distances of each of the levels of the MSF as shown in Section III.2. Moreover, an intentional oversupply or undersupply of resource states between the zones can be used to tune various space–time trade-offs for FTQC. For example, a magic state factory (MSF) may require multiple magic state distillation units to keep the core processor fully stacked, especially when a reaction-time-limited compilation is used to expedite the consumption rate of magic states [7]. A smaller number of distillation units may still be able to fault-tolerantly execute the quantum algorithm if some slowdown caused by occasionally empty buffer registers can be tolerated. However, reducing the (logical) size of the MSF increases the idle time in the core processor, which increases the overall accumulated error rates, potentially requiring larger code distances. Therefore it is not a priori obvious that such a strategy will allow for longer runtime on a smaller quantum computer without taking the noise profile of the hardware into account. Our study provides a detailed approach to overcoming the challenges of analyzing such trade-offs.
We model the optimization of space and time as a bi-objective optimization problem subject to a total error budget constraint discussed in Section III and Appendix B. The problem is solved using a simple and effective heuristic that determines the sizes of all levels of the MSF and the code distances required by all logical qubits in the MSF and the core processor, while meeting a total accumulated error budget. By solving this optimization problem, we observe that the space and time costs of any FTQC can be efficiently predicted from a small set of key attributes of (a) the quantum program, and (b) the quantum computer executing it. Firstly, the two relevant attributes of the FTQC program are denoted by the Greek small letters $\alpha$ and $\beta$, representing the average size of lattice surgeries in the core processor relative to the number of logical qubits and a user-defined slowdown factor within the buffers, respectively. Secondly, the noise profile of the QPU is also important for the assembler. We observe that this information can also be reduced to a small number of parameters for a predictive model of the logical error rates of each FTQC protocol, denoted by $\mu$ and $\Lambda$, respectively representing the logical error prefactor and the logical error suppression rate of the protocol. Prior literature have confined to using the quantum memory on a single surface code patch, $\Lambda=\Lambda_{\text{mem}}$, to approximate the performance of all lattice surgeries in the FTQC program [8]. However, a more comprehensive predictive model must incorporate the exponential logical error suppression factors of all fault-tolerant protocols in the assembly. Note that the quantum memory error suppression rate $\Lambda_{\text{mem}}$ describes the asymptotic multiple of improvement in the logical fidelity of the QEC code when increasing its distance by 2. Therefore, in Appendix B we develop more general logical error rate models for lattice surgeries with various shapes and involving various numbers of surface code patches.
The paper is organized as follows. Section II introduces our proposed architecture. Section III formalizes the bi-objective optimization problem to be solved when assembling the proposed architecture and the optimization framework we used to solve the problem. In Section IV we present several applications of our systematic assembly procedure by numerically studying FTQC execution space–time trade-offs with respect to specifications of the input quantum circuits, the quantum hardware noise profile, and the reaction time of classical decoders. Lastly, our concluding remarks follow in Section V.
@@ II The Fault-Tolerant Architecture
We assume rotated surface codes as our QEC scheme of choice and assume that the program to be run is transpiled to multi-qubit Pauli rotations of the angles $\pi$, $\pi/2$, and $\pi/4$ for Clifford operations, and $\pi/8$ for non-Clifford operations [9, 10]. These Pauli rotations are in turn implemented using fault-tolerant lattice surgery [5, 6, 4]. A core processor is designed to sequentially perform multi-qubit, long-range entanglements for sets of Pauli rotation measurements using auxiliary qubits to connect all the logical qubits required by a given logical operation. Clifford operations can be implemented via lattice surgery using the logical qubits involved with an auxiliary qubit initialized in the |$\ket{0}$ state, while non-Clifford operations require entangling the logical qubits with a qubit in a magic state [5]. The core (see Section II.1) is connected to an external multi-level MSF, as detailed in Section II.2, that supplies high-fidelity magic states required for the implementation of the non-Clifford logical operations.
@@ II.1 The Core Processor
The core processor is the central component of the architecture, where logical operations are performed on the logical qubits. It comprises two key elements: the memory fabric and a buffer. The memory fabric is responsible for storing the quantum information required by the computation and performing its logical operations, while the buffer is used to simplify access to magic states so that the memory fabric does not have to directly interface with the MSF. There, the resources required by the $\pi/8$ rotations are prepared and held until they are consumed.
@@
An example layout of the core processor is shown in Fig. 1  for a circuit requiring 18 logical qubits to store its data. We follow the fast block layout of Ref. [5] for the memory fabric, where logical qubits are arranged in two-tile, two-qubit data patches connected to a quantum bus using a square arrangement for performing lattice surgery. This is a time-efficient layout as all Pauli operators—$X$, $Z$, and $Y$ (the combination of both)—are directly accessible by the quantum bus. In an exact square arrangement, the $Q$ logical qubits requested in the algorithm would require a space of $12Q+\sqrt{8Q}+$ 1 tiles plus the tiles required for the buffer space. To save on space, we shorten the last column when $\sqrt{Q/2}$ is not an integer. Different core layouts, such as the compact block layout of Ref. [5] and the layout of Ref. [12] for edge-disjoint path compilation, can be incorporated into our models by adjusting the core processor size accordingly. However, a significant drawback of these layouts is their inability to provide access to the $Y$ operator of data qubits within a single logical cycle, which commonly appear in circuits after applying the transpilation procedure described in Refs [4] and [5]. This limitation requires converting logical gates with $Y$ operators into gates with only $X$ and $Z$ operations and commuting the additional Clifford gates out to avoid additional overhead in terms of time.
@@
Figure 1: Example layout of the core processor. The memory fabric is composed of $Q=18$ logical qubits distributed in nine two-tile, two-qubit data patches surrounded by bus patches, while the buffer is composed of patches dedicated to storing magic states produced in the MSF. This storage space is designed to allow post-corrected $\pi/8$ rotations by connecting the magic state growth patches to a correction qubit patch using lattice surgery in parallel to the lattice surgery used to connect the data patches to the magic state consumed in a magic state storage patch. Magic state growth patches also allow magic state expansion if the produced magic states are of a different size than is required. The sides of the buffer are used to store correction qubits. The number of correction storage patches is determined by the time required to perform the classical processing in the post-corrected protocol.
The buffer in Fig. 1 is designed to efficiently implement the post-corrected $\pi/8$ rotation protocol described in Ref. [5] (see Appendix A for more details). Within this buffer, magic states produced by the MSF are sent to magic state growth patches where they are enlarged to match the code distance used in the core processor. Next, the correction state is created in the correction patch following the post-corrected protocol. The magic state used in the correction state creation is then moved to a magic state storage patch using patch deformation, while the correction qubit is moved to a correction storage patch, freeing their spaces for the next operation. Once stored, the magic state is ready for consumption by a $\pi/8$ rotation. To apply the rotation, the magic state is entangled with the data qubits in the memory fabric required for the operation using lattice surgery. After that, the $X$-basis single-qubit measurement of the post-corrected protocol is performed in the patch that was holding the magic state. At the earliest available time, the correction state entangled to it is measured in the $X/Z$-basis, determining whether a Pauli correction is required to define whether the rotation should have been a $+\pi/8$ or $-\pi/8$ rotation. This Pauli correction is not executed physically as it can be commuted to the end of the circuit, updating the axis of subsequent rotations based on whether they anticommute.
The time required to classically process these measurement results and determine subsequent operations is referred as the reaction time. The reaction time determines whether $\pi/8$ rotation computation is $T$-count-limited or reaction-time-limited. In $T$-count-limited computation, each $\pi/8$ rotation is executed sequentially (or in parallel if subsequent rotations commute following a compilation procedure [4]). In reaction-time-limited computation, quantum teleportation can be used to parallelize non-commutable $\pi/8$ rotations when the reaction time is faster than the time needed to complete a logical cycle (see Appendix A for further details).
The number of correction storage patches in the buffer depends on the reaction time, as a patch remains occupied until this classical processing has completed. The buffer is designed to allow the simultaneous preparation of magic and correction states without introducing delays. If the MSF produces magic states sufficiently fast to keep the buffer continuously active, a magic state will always be available for consumption at the start of every logical cycle.
Additional buffers can be placed around the memory fabric to enable the parallel preparation of multiple magic states, supporting parallel $\pi/8$ rotations. If the MSF production rate is slower than one magic state per logical cycle, the buffer size can be adjusted accordingly to minimize space requirements.
@@ II.2 Magic State Factory
The MSF is a module of the fault-tolerant architecture dedicated to the production of high-fidelity magic states required to implement the $\pi/8$ rotation gates. Magic states are initially prepared using physical operations in a magic state preparation unit. As these are highly error-prone states that most likely have fidelities below those required for fault tolerance, rounds of magic state distillation are performed using magic state distillation units, where the magic states at one particular level are consumed to create higher-fidelity magic states. Once ready, the magic states distilled are transferred to a distillation unit in the next level where they are expanded to a larger code distance, if required, in magic state growth units. When magic states reach the required fidelity, the distillation process has been completed and the magic states created at the highest distillation level are sent to the core processor to be consumed when performing the $\pi/8$ rotation gates. Figure 2 shows an example of the architecture described.
Figure 2: Architecture of the MSF and its interface with the core processor buffer. Magic states are prepared using dedicated preparation units following a preparation protocol. Once preparation succeeds, magic states wait until a storage patch is available and then are moved to there using patch deformation. When in a storage patch, a similar post-correction protocol as is performed in the buffer follows, where a correction qubit is created, entangled with the magic state, and stored. The magic state is then used for the $\pi/8$ rotations performed between this magic state and data and distilling port qubits within the distillation unit. The higher-fidelity magic states are prepared in the distilling ports of these units. Once ready, they proceed to growth units dedicated to a next-level distillation unit. The process repeats until the highest level sends a magic state to the core processor. In this example, two distillation levels are represented for 15:1 distillation protocols, composed of three and two distillation units each from lowest to highest. While the lower-level unit design allows distillation of higher-fidelity magic states every 13 logical cycles if enough preparation units are available, topological limitations make the higher-level unit distill magic states only every 15 logical cycles instead due to the extra time required to load magic states into the data qubits before a distillation cycle begins.
@@
As a good example that is used as a testbed throughout this paper, the 15:1 distillation protocol [13] uses 15 lower-fidelity magic states and produces a single higher-fidelity magic state. An analysis of this distillation protocol shows that if the input error rates of the magic states are $e_{\text{in}}$ and the logical error rates of the individual logical gates are $e_{\text{mem}}$, then the output error is bounded by $O(e_{\text{in}}^{3}+e_{\text{mem}})$ from the distillation protocol. From this, one can determine the output fidelity of a single magic state distillation unit given the input fidelities and code distances. Additionally, it is possible to approximate the probability of a successful distillation for this protocol using numerical simulations [6]. Figure 2 shows an architecture for the 15:1 protocol. The first-level distillation units are customized to run in 13 logical operations, accounting for 11 gates, one cycle for correction, and one cycle for emptying the distillation port, at a given code distance. Higher-level distillation units require two additional logical steps to load the four data (or stabilizer) qubits before starting a new distillation cycle due to topological constraints. This architecture requires at least 28 surface code tiles per distillation unit, even without accounting for the correction storage and magic state preparation space, which is much greater than the five logical qubits required for the protocol itself.
@@
The very first level of the multi-level MSF consumes logical magic states prepared from physical non-Clifford states that are fault-tolerantly grown to the logical state of a full surface code patch. These protocols generally involve the creation of a physically relevant state (such as a physical magic state), and then the injection of this magic state into a larger quantum error-correcting code [16, 17]. The result is a magic state of fidelity $e_{\text{prep}}$ at a specified code distance. Different protocols require different physical operations, and have different time scales, but most require $\sim$1–2 logical cycles, where a logical cycle requires $d$ rounds of QEC at the final code distance $d$. The physical execution of the protocols on a 2D layout causes additional error to be accumulated which can also be numerically estimated. At the end of the process, we obtain the error rate $e_{\text{in}}$ for the initial level of magic state distillation. These preparation protocols generally have a probability of failure, and these probabilities can be numerically estimated.
@@
Although distillation levels can be encoded using the same code distances, it is more space efficient to use different encodings for the logical qubits at different levels, usually in increasing order. Therefore, a growth protocol must be used between the MSF levels to increase the code distance of the magic states to match the distance of the next distillation level or that of the core processor. These growth protocols operate similarly to the preparation protocols, in which a state encoded in a smaller state is surrounded by additional qubits in specific states, followed by QEC rounds in the grown patch. These growth protocols also have a small probability of failure that can be numerically estimated.
@@
Magic state distillation is a probabilistic procedure, and thus each distillation unit has some chance of failing to output a state. To reduce the probability of no magic state existing when one is needed in our proposed architecture, we can set additional distillation units to ensure the system operates in a steady state, ensuring the average production is above consumption. This can be complemented with the inclusion of a buffer register between levels where magic states can remain idle while waiting to be consumed without blocking the start of the next distillation cycle in the unit where it was produced. These zones essentially need only to be a single logical tile where the magic state can sit, but having access to the tiles to both load and unload the magic states means they should generally be lined up.
III The Assembly Process